# Optimizing Allreduce Operations for Heterogeneous Architectures with Multiple Processes per GPU

### v1

@@
Abstract.
Large inter-GPU all-reduce operations, prevalent throughout deep learning, are bottlenecked by communication costs. Emerging heterogeneous architectures are comprised of complex nodes, often containing $4$ GPUs and dozens to hundreds of CPU cores per node. Parallel applications are typically accelerated on the available GPUs, using only a single CPU core per GPU while the remaining cores sit idle. This paper presents novel optimizations to large GPU-aware all-reduce operations, extending lane-aware reductions to the GPUs, and notably using multiple CPU cores per GPU to accelerate these operations. These multi-CPU-accelerated GPU-aware lane all-reduces yield speedup of up to 2.452.45x for large MPI all-reduces across the NVIDIA A100 GPUs of NCSA’s Delta supercomputer. Finally, the approach is extended to NVIDIA’s and AMD’s collective communication libraries, achieving speedup of up to 1.771.77x and 1.711.71x, respectively, across $2$ state-of-the-art supercomputers.
Allreduce, heterogeneous, GPU, MPI, NCCL, RCCL, multi-lane, node-aware
$c$s: Computing methodologies Parallel algorithmsccs: Computing methodologies Massively parallel algorithmsccs: Computing methodologies Distributed algorithmsccs: Computer systems organization Parallel architectures
1. Introduction
@@
Figure 1. Example All-Reduce Setup
[…]
Example buffers to be reduced across $8$ GPUs and $2$ nodes.
@@
While there are many existing all-reduce algorithms, the standard approach for large data sizes is through a ring pattern (ring0, ), in which every process exchanges a portion of the data with neighbors at each step, the first of which is displayed in Figure 2.
Figure 2. Ring All-Reduce, First Step
[…]
The first step of a ring all-reduce.
The cost of each message varies with the relative locations of the sending and receiving processes (perf_model_locality, ), and the ring algorithm attempts to optimize locality by restricting communication to neighboring ranks. Assuming an optimal process layout, the majority of processes may communicate locally, such as within a node. For example, the majority of messages displayed in Figure 2 are intra-node. However, at the boundary, processes must communicate with neighboring nodes at each step.
The multi-lane all-reduce (lane, ) improves the cost of large reductions through lane-awareness, with each process per node performing a separate portion of the inter-node all-reduce, exemplified in Figure 3.
Figure 3. Lane All-Reduce, Inter-Node Step
[…]
The inter-node step of a multi-lane all-reduce.
@@
This paper explores methods for optimizing large GPU-aware all-reduce operations for emerging heterogeneous architectures. This paper presents a novel extension of the multi-lane algorithm to GPU-aware all-reduces. The authors further optimize the all-reduce through improved resource utilization, with multiple CPU cores each progressing a portion of the all-reduce asynchronously. Finally, the authors apply an equivalent technique to the NVIDIA and AMD collective communication libraries (CCLs) (NVIDIA2025NCCL, ; AMD2025RCCL, ), using multiple processes per GPU to accelerate the widely used CCL all-reduce.
Figure 4. Standard MPI_Allreduce vs our optimized all-reduce
[…]
@@
  • Compares the performance of GPU-aware standard and multi-lane algorithms;
@@
  • Presents a performance analysis of these optimizations, achieving up to 2.452.45x speedup over OpenMPI’s GPU-aware all-reduce; and
  • Accelerates NCCL and RCCL all-reduce operations by up to 1.771.77x and 1.711.71x, respectively, through the use of multiple processes per GPU.
The remainder of our paper is outlined as follows. Section 2 describes background information and related works, detailing the baseline ring and multi-lane all-reduce algorithms. The methodology for optimizing large all-reduces is described in Section 3. Section 4 provides performance results for our approaches, with MPI results shown in Section 4.1 and extensions to the CCLs shown in Section 4.2. Finally, Section 5 provides concluding remarks.
2. Background
@@
All-reduce operations are generally performed using three different algorithms: recursive doubling (recursive_doubling, ), Rabenseifner’s algorithm (ring0, ), and ring (ring0, ). Recursive doubling is used for small reductions, minimizing the number of messages to $\log{p}$ where $p$ is the number of processes, but sending all $n$ bytes of the buffer in each message. Rabenseifner’s algorithm improves performance for moderately-sized buffers, with each process reducing a fraction of the buffer before gathering all reductions. The total number of messages increases to $2\cdot\log{p}$, but the amount of data communicated and reduced upon is cut in half at each step. The ring all-reduce further optimizes for large reductions by restricting communication only to immediate neighbors, increasing the message count to $2\cdot(p-1)$ (ring0, ; ring1, ). In recent years, this algorithm has been adopted within deep learning applications (ring2, ; ring3, ).
@@
Input: Rank, count, size from group communicator
@@
Multi-lane collectives (lane, ) reduce the variability between message costs through node-awareness. The multi-lane all-reduce, detailed in Algorithm 2, first performs a reduce-scatter on-node so that each process within the node holds a fraction of the buffer. Then, each process performs an all-reduce with its corresponding process on each other node. Finally, the reduced buffer portions are gathered among all processes within a node. Each process sending an equal portion of the buffer allows for the operation to better reach injection bandwidth limits since per-process inter-node messages are reduced in both size and count. Multi-lane algorithms have shown to greatly improve the performance of collectives on many-core architectures and heterogeneous systems requiring copies of the device buffer to host memory (gpu_c2c_lane, ). This paper presents a novel extension and analysis of these algorithms to GPU-aware all-reduce operations on heterogeneous systems using remote direct memory access (RDMA) communication.
2.1. GPU-Aware All-Reduce Operations
GPU-aware MPI implementations allow the CPU to offload communication to the GPU and network interface card (NIC) by utilizing the GPUDirect RDMA protocol. The NICs on emerging supercomputers support RDMA, allowing data to flow directly between the NIC and the GPU, bypassing copies with host memory. Large GPU-aware all-reduce operations take advantage of GPUDirect RDMA. While manually copying data to the CPU showed benefits on Power9 systems (perf_hetero, ), more modern systems obtain the best performance through the use of GPUDirect RDMA communication.
@@ 2.1. GPU-Aware All-Reduce Operations
On many-core CPU systems, similar synchronization overheads have been reduced through partitioned communication, which allows each thread to instantiate communication on a local portion of the buffer without waiting for all threads to finish computation (partitioned, ). Similarly, partitioned communication allows for portions of the buffer to be received early, so that each thread can begin local computation as soon as its portion of the receive arrives, rather than waiting for the full buffer. While partitioned communication has been shown to optimize many operations on CPU-only systems, there have been extensions to heterogeneous systems. Due to the high overheads of kernel launches, along with the inability to initiate communication directly from the GPU, partitioning heterogeneous data into smaller partitions can incur bottlenecks.
@@
2.2. The CCLs
Proprietary communication libraries, such as NVIDIA’s Collective Communications Library (NCCL) and AMD’s ROCm Communication Collectives Library (RCCL), implement the all-reduce with increased knowledge about the GPU and NIC hardware in use, yielding optimizations over MPI implementations that make them commonplace in AI applications. These all-reduce operations achieve very high bandwidth, optimal for large reductions. Similar to MPI, these all-reduces are typically performed with a single process per GPU. This paper presents a novel optimization to CCL all-reduces in which multiple processes per GPU each initiate a separate CCL all-reduce on a unique subset of the buffer.
2.3. Related Work
@@
Other collectives such as the all-to-all, all-gather and related neighborhood methods feature a similar communication pattern to the all-reduce where in the worst case, every pair of processes must communicate. Intra-node communication has been optimized with shared memory (shared_memory, ; GrahamSharedMemoryColl2008, ; JainSharedMemoryColl2018, ) on CPUs. Hierarchical collectives aggregate all data onto a single leader per node, reducing inter-node traffic (hierarchical, ; KaronisHierColl2000, ; GrahamHierCheetah2011, ; TraffHierAllgather2006, ). Multi-leader collectives perform hierarchical algorithms with multiple leaders per node (multileader_hierarchical, ; multileader_hierarchical2, ). Locality-aware aggregation has also been used to aggregate messages, reducing the number of inter-node steps, both in the context of small all-reduce operations (bienz_allreduce, ), as well as other collective operations (bienz_bruck_allgather, ; locality_aware_alltoall, ).
@@
Topology-aware approaches for collective operations have also been developed. For instance, for a cluster layout consisting of a two-dimensional grid of nodes where each node is directly connected to its four neighboring nodes, an all-reduce can be performed twice, once in each axis direction (BhateleTorusColl2012, ; torus, ), reducing the number of messages from the total number of nodes to the sum of all axes dimensions. Topology-aware mapping is commonly used to map processes within other topologies (topo0, ; topo1, ; topo2, ), such as trees (PatarasukTreeAllreduce2007, ), to minimize long-distance messages.
@@
GPU-aware optimizations for collectives have been widely explored in recent years. Hierarchical methods have been analyzed on heterogeneous architectures (infini_hierarchical, ; gpu_multileader_hierarchical, ; hierarchical_ccl, ), communication primitives and collectives have been optimized to use direct memory access between GPUs more efficiently (gpu_dma, ; gpu_dma2, ; gpu_dma3, ; gpu_dma4, ), and compressed collectives have been used to reduce communication costs (comp_cosmo, ; comp_alltoall, ; comp_allreduce, ; zccl, ).
3. Methods
((a)) 8 nodes, varying buffer sizes
((b)) $2^{27}$ floats, varying node counts
Figure 5. Multi-Lane GPU-Aware Performance
@@ 3. Methods
Baseline GPU-aware all-reduce operations implement standard algorithms, such as the ring and multi-lane implementations in Algorithms 1 and 2, using GPUDirect RDMA for all communication. Further, these baseline implementations use a single MPI process per GPU.
3.1. Multiple Processes Per GPU
The baseline all-reduce operations are further optimized with multiple processes per GPU.
Figure 6. Partitioning Buffer Over Multiple Processes Per GPU
[…]
@@ 3.1. Multiple Processes Per GPU
Each GPU on a node is assigned to an equal number of unique processes, as exemplified in Figure 6. In this example, there are two processes per GPU, displayed as red and blue outlines. The buffers to be reduced are partitioned evenly across the multiple processes per node.
@@
3.1.1. Standard Approach
@@
Figure 7. Multi-PPG, Multi-Lane
[…]
The number of kernel launches per process remains constant. As the MPI all-reduce operation does not allow passing streams and the NVIDIA Multi-Process Service (MPS) is assumed to be disabled, the local reductions are unable to be performed in parallel. However, as there is asynchrony among processes, all kernels are unlikely to compete for the GPU at one time.
3.1.2. Multi-lane Approach
@@
4. Results
((a)) 8 nodes, varying buffer sizes
((b)) $2^{27}$ floats, varying node counts
Figure 8. Standard and multi-lane (with MPI_Allreduce and ring all-reduce) algorithms with the maximum of multiple-processes per GPU
The impact of multi-lane and multi-process per GPU all-reduce algorithms is analyzed against OpenMPI, NCCL, and RCCL on state-of-the-art supercomputers: NCSA’s Delta and DeltaAI as well as LLNL’s Tuolumne. Some details of these systems are provided below.
Delta: This system provides a variety of nodes, including both NVIDIA and AMD GPUs. However, due to the limited number of AMD nodes, only NVIDIA nodes were considered for the scaling studies in this paper. All-reduce operations were analyzed on the single-socket nodes, each composed of a 64 core AMD Milan CPU with 4 NVIDIA A100 GPUs. DeltaAI: This supercomputer is comprised of Grace Hopper superchips. Each node is equipped with four NVIDIA GH200 superchips, each containing a Grace ARM 72-core CPU and a NVIDIA H100 GPU. Tuolumne: Each node has four AMD MI300A APUs, each of which has a 24-core AMD EPYC CPU and a CDNA 3 GPU.
The impact on MPI all-reduce operations is analyzed solely on Delta, as OpenMPI was available only on this system. All other systems provide Cray MPICH, which forbids any buffer from being communicated through GPU-aware MPI operations if it was not allocated by the calling process. As a result, sharing IPC buffers within GPU-aware MPI is not possible on these systems. The impact of multiple processes per GPU on the CCL libraries, however, was evaluated on Cray systems.
Note that NVIDIA MPS was not enabled during benchmarking.
4.1. MPI Analysis
All MPI all-reduce tests were analyzed using OpenMPI 5.0.5 and CUDA 12.4.0. All processes were assigned to the GPU adjacent to their NUMA region. Each benchmark was run $10$ separate times to demonstrate reproducibility. All plots contain error bars showing the run to run variation of the algorithms.
Figure 5 compares the performance of a standard MPI_Allreduce operation against the multi-lane optimization. These tests are using a single process per GPU, and using all four GPUs available per node on Delta. While the standard algorithm outperforms multi-lane at the smallest data sizes, the multi-lane algorithm yields significant speedups for larger all-reduces. The speedups for the multi-lane algorithm increase with node count, achieving 2.12.1x speedups at $32$ nodes.
Figure 9. A standard MPI_Allreduce with multiple processes per GPU.

### v2

# Optimizing Allreduce Operations for Modern Heterogeneous Architectures with Multiple Processes per GPU

@@
Abstract
Large inter-GPU all-reduce operations, prevalent throughout deep learning, are bottlenecked by communication costs. Emerging heterogeneous architectures are comprised of complex nodes, often containing $4$ GPUs and dozens to hundreds of CPU cores per node. Parallel applications are typically accelerated on the available GPUs, using only a single CPU core per GPU while the remaining cores sit idle. This paper presents novel optimizations to large GPU-aware all-reduce operations by extending the lane-aware algorithm to heterogeneous architectures and notably using multiple CPU cores per GPU to accelerate these operations. Using GPUDirect RDMA and host copy communications respectively, these multi-CPU-accelerated GPU-aware all-reduces yield speedups over system MPI of up to 33x on LLNL’s Tuolumne supercomputer and up to 2.452.45x for large MPI all-reduces across the NVIDIA A100 GPUs of NCSA’s Delta supercomputer.
I Introduction
@@
Figure 1: Example All-Reduce Setup
@@
While there are many existing all-reduce algorithms, the standard approach for large data sizes is through a ring pattern [35], in which every process exchanges a portion of the data with neighbors at each step, the first of which is displayed in Figure 2.
Figure 2: Ring All-Reduce, First Step
The cost of each message varies with the relative locations of the sending and receiving processes [7], and the ring algorithm attempts to optimize locality by restricting communication to neighboring ranks. Assuming an optimal process layout, the majority of processes may communicate locally, such as within a node. For example, the majority of messages displayed in Figure 2 are intra-node. However, at the boundary, processes must communicate with neighboring nodes at each step.
The multi-lane all-reduce [40] improves the cost of large reductions through lane-awareness, with each process per node performing a separate portion of the inter-node all-reduce, exemplified in Figure 3.
Figure 3: Lane All-Reduce, Inter-Node Step
@@
This paper explores methods for optimizing large GPU-aware all-reduce operations for emerging heterogeneous architectures. This paper presents a novel extension of the multi-lane algorithm to GPU-aware all-reduces backed by host copy and GPUDirect Remote Direct Memory Access (RDMA) communication on NVIDIA Ampere-based and AMD MI300A-based systems, respectively. We further optimize the all-reduce through improved resource utilization, with multiple CPU cores each progressing a portion of the all-reduce asynchronously with performance as seen in Figure 4. Finally, we analyze the performance of the all-reduces in the various partitioning modes of the AMD MI300A, showing that the multi-lane all-reduces benefit from increased logical GPU count.
Figure 4: Standard MPI_Allreduce vs our optimized all-reduce on LLNL’s Tuolumne
@@
  • Compares the performance of GPU-aware standard and multi-lane algorithms on modern architectures;
@@
  • Presents a performance analysis of these optimizations on modern architectures, achieving up to 2.452.45x speedup over OpenMPI’s GPU-aware all-reduce using host copy communication and up to 1.171.17x speedup over Cray MPICH’s GPUDirect RDMA implementation; and
  • Accelerates all-reduces using the TPX and CPX partitioning modes using the multi-lane algorithm, achieving up to 33x speedup over the Cray MPICH implementation.
The remainder of our paper is outlined as follows. Section II describes background information and related works, detailing the baseline ring and multi-lane all-reduce algorithms. Section III provides benchmarks of point-to-point communication using multiple processes per GPU to motivate our work. The methodology for optimizing large all-reduces is described in Section IV. Section V provides performance results for our approaches, with results using host copy communication shown in Section V-A and extensions to GPUDirect RDMA communication on the AMD MI300A shown in Section V-B. Performance on the various partitioning modes of the AMD MI300A is shown in Section V-B1. Finally, Section VI provides concluding remarks.
II Background
All-reduce operations are generally performed using three different algorithms: recursive doubling [38], Rabenseifner’s algorithm [35], and ring [35]. Recursive doubling is used for small reductions, minimizing the number of messages to $\log{p}$ where $p$ is the number of processes, but sending all $n$ bytes of the buffer in each message. Rabenseifner’s algorithm improves performance for moderately-sized buffers, with each process reducing a fraction of the buffer before gathering all reductions. The total number of messages increases to $2\cdot\log{p}$, but the amount of data communicated and reduced upon is cut in half at each step. The ring all-reduce further optimizes for large reductions by restricting communication only to immediate neighbors, increasing the message count to $2\cdot(p-1)$ [35, 33]. In recent years, this algorithm has been adopted within deep learning applications [15, 37].
@@ II Background
Input: rank, count,
   size from group communicator
@@
Multi-lane collectives [40] reduce the variability between message costs through node-awareness. The multi-lane all-reduce, detailed in Algorithm 2, first performs a reduce-scatter on-node so that each process within the node holds a fraction of the buffer. Then, each process performs an all-reduce with its corresponding process on each other node. Finally, the reduced buffer portions are gathered among all processes within a node. Each process sending an equal portion of the buffer allows for the operation to better reach injection bandwidth limits since per-process inter-node messages are reduced in both size and count. Multi-lane algorithms have shown to greatly improve the performance of collectives on older many-core architectures and heterogeneous systems requiring copies of the device buffer to host memory [13]. This paper presents a novel extension and analysis of these algorithms to GPU-aware all-reduce operations on heterogeneous systems using GPUDirect RDMA communication; this paper also justifies their continued applicability on modern systems requiring copies between host and device memory.
II-A GPU-Aware All-Reduce Operations
GPU-aware MPI implementations allow the CPU to offload communication to the GPU and network interface card (NIC) by utilizing the GPUDirect RDMA protocol. The NICs on emerging supercomputers support RDMA, allowing data to flow directly between the NIC and the GPU, bypassing copies with host memory. Large GPU-aware all-reduce operations take advantage of GPUDirect RDMA. While manually copying data to the CPU showed benefits on Power9 systems [9], more modern systems obtain the best performance through the use of GPUDirect RDMA communication.
@@ II-A GPU-Aware All-Reduce Operations
On many-core CPU systems, similar synchronization overheads have been reduced through partitioned communication, which allows each thread to start communication on a local portion of the buffer without waiting for all threads to finish computation [19]. Similarly, partitioned communication allows for portions of the buffer to be received early, so that each thread can begin local computation as soon as its portion of the receive arrives, rather than waiting for the full buffer. While partitioned communication has been shown to optimize many operations on CPU-only systems, there have been extensions to heterogeneous systems [13]. Due to the high overheads of kernel launches, along with the inability to initiate communication directly from the GPU, partitioning heterogeneous data into smaller partitions can incur bottlenecks.
(a) (Delta)
(b) Default SPX mode (Tuolumne)
(c) Fine-grained CPX mode (Tuolumne)
Figure 5: A ping-pong benchmark using multiple active processes per (logical) GPU featuring performance on Delta (left), and the SPX (middle) and CPX (right) modes of the AMD MI300A on Tuolumne
@@
II-B Partitioning Modes of the AMD MI300A
The AMD MI300A Accelerated Processing Unit (APU) is a processor that places the CPU and GPU in a single package [1]. There are three CPU chiplets each with 8-core AMD Zen4 CPUs and six Accelerator Complex Dies (XCDs) with 38 CDNA3-based GPU compute units each. 128 GB of HBM3 memory is accessible to both CPU cores and GPU compute units backed by 256 MB of shared last level cache.
At runtime, the XCDs can be grouped logically depending on the user’s needs. By default, the MI300A APU operates in Single Partition X-celerator (SPX) mode in which all 6 XCDs are exposed to the user as a single GPU. In this case, the workload on the GPU is automatically distributed across all XCDs. With the Triple Partition X-celerator (TPX) mode, the APU is treated as three GPUs of two XCDs each. In Core Partitioned X-celerator (CPX) mode, each XCD is exposed as a single GPU.
The TPX and CPX partitioning modes offer the developer more granular control over the hardware that is chosen to execute their applications. Furthermore, since all partitioning modes can access the entire pool of HBM3 memory, these modes offer increased parallelism when used effectively.
II-C Related Work
@@
Other collectives such as the all-to-all, all-gather and related neighborhood methods feature a similar communication pattern to the all-reduce where in the worst case, every pair of processes must communicate. Intra-node communication has been optimized with shared memory [20, 17, 23] on CPUs. Hierarchical collectives aggregate all data onto a single leader per node, reducing inter-node traffic [45, 27, 18, 41]. Multi-leader collectives perform hierarchical algorithms with multiple leaders per node [26, 4]. Locality-aware aggregation has also been used to aggregate messages, reducing the number of inter-node steps, both in the context of small all-reduce operations [8], as well as other collective operations [6, 10].
@@
Topology-aware approaches for collective operations have also been developed. For instance, for a cluster layout consisting of a two-dimensional grid of nodes where each node is directly connected to its four neighboring nodes, an all-reduce can be performed twice, once in each axis direction [5, 29], reducing the number of messages from the total number of nodes to the sum of all axes dimensions. Topology-aware mapping is commonly used to map processes within other topologies [12, 30, 28], such as trees [32], to minimize long-distance messages.
@@
GPU-aware optimizations for collectives have been widely explored in recent years. Hierarchical methods have been analyzed on heterogeneous architectures [16, 39, 21], communication primitives and collectives have been optimized to use direct memory access between GPUs more efficiently [24, 34, 42, 11], and compressed collectives have been used to reduce communication costs [25, 43, 44, 22].
Proprietary communication libraries, such as NVIDIA’s Collective Communications Library (NCCL) [31] and AMD’s ROCm Communication Collectives Library (RCCL) [2], implement the all-reduce with increased knowledge about the GPU and NIC hardware in use, yielding optimizations over MPI implementations that make them commonplace in AI applications. These all-reduce operations achieve very high bandwidth, optimal for large reductions. Similar to MPI, these all-reduces are typically performed with a single process per GPU.
III Microbenchmarks
GPU-aware operations, such as the all-reduce, are performed using one controlling process per GPU. This process initiates all GPU-aware communication involving its GPU. However, modern systems feature many more CPU cores than GPUs, causing them to be underutilized.
To demonstrate the potential advantage of using multiple processes per GPU, we evaluated a ping-pong benchmark on both systems analyzed. In this benchmark, there are two nodes where each rank on a node sends and receives a message with its corresponding rank on the other node. Each GPU on a node is participating in the benchmark and controlled by at least 1 MPI process. To obtain the time it takes to send or receive one message, the round-trip time is divided by $2$.
Benchmark results on the Delta system with intermediate copies between host and device memory can be seen in Figure 5(a). When 16 processes per GPU are used, this yields a significant speedup of 3.493.49x over a standard ping-pong of the largest message size, suggesting a benefit by the use of multiple processes per GPU during communication on Delta.
Results using GPUDirect RDMA communication on the SPX and CPX partitioning modes of the APU on Tuolumne can be seen in Figures 5(b) and 5(c). Similarly to results on Delta, the ping-pong using SPX mode yields speedups of up to 1.891.89x over a standard ping-pong when using 16 processes per GPU. However, when fine-graned control of each XCD is enabled with CPX mode, using multiple processes per GPU does not show significant benefit. Therefore, these results suggest using mutiple processes per GPU is more applicable when using the coarse-grained partitioning on Tuolumne’s APUs.
IV Methods
@@
Baseline GPU-aware all-reduce operations implement standard algorithms, such as the ring and multi-lane implementations in Algorithms 1 and 2. Further, these baseline implementations use a single MPI process per GPU.
IV-A Multiple Processes Per GPU
@@
The baseline all-reduce operations are further optimized with multiple processes per GPU. Each GPU on a node is assigned to an equal number of unique processes, as exemplified in Figure 6. In this example, there are two processes per GPU, displayed as red and blue outlines. The buffers to be reduced are partitioned evenly across the multiple processes per node.
@@
IV-A1 Standard Approach
Figure 6: Partitioning Buffer Over Multiple Processes Per GPU
@@ IV-A1 Standard Approach
The number of kernel launches per process remains constant. The host copy implementations are likely to be performed in parallel excluding the host-device copy operations since the Multi-Process Service (MPS) is disabled on the Ampere-based system. For the GPUDirect RDMA implementations, communication is either parallelized using MPS or, as there is asynchrony among processes, all kernels are unlikely to compete for the GPU at one time. It is noted on the AMD MI300A system tested, AMD has natively supported MPS features in their GPU driver [14], so MPS is implied to be enabled.
IV-A2 Multi-lane Approach
@@
V Results
The impact of multi-lane and multi-process per GPU all-reduce algorithms is analyzed against OpenMPI and Cray MPICH on state-of-the-art supercomputers: NCSA’s Delta and LLNL’s Tuolumne. Some details of these systems are provided below.
Delta: This system provides a variety of nodes, including both NVIDIA and AMD GPUs. The number of AMD nodes is limited, so only NVIDIA nodes were benchmarked for this paper. All-reduce operations were analyzed on the single-socket nodes, each composed of a 64 core AMD Milan CPU with 4 NVIDIA A100 GPUs.
Tuolumne: Each node has four AMD MI300A APUs, each of which has a 24-core AMD EPYC CPU and a CDNA 3 GPU.
The impact on MPI all-reduce operations using copies between host and device memory is analyzed solely on Delta, as only OpenMPI was available on this system at the time of writing. Tuolumne supports GPUDirect RDMA provided by Cray MPICH. Cray MPICH does not support any buffer from being communicated through GPU-aware MPI operations if it was not allocated by the calling process. As a result, sharing IPC buffers within GPU-aware MPI is not possible on these systems. Thus, each reduction using multiple processes per GPU analyzed on Tuolumne copies to and from another device buffer allocated by the calling process before calling the MPI routines on the local buffer.
Note that NVIDIA MPS was not enabled during benchmarking on Delta whereas an equivalent implementation by AMD is implicitly enabled on Tuolumne.
Figure 8: Multi-lane GPU-aware performance on 8 nodes of Delta, varying buffer sizes
V-A Host Copy Analysis
All all-reduce tests on Delta were analyzed using OpenMPI 5.0.5 and CUDA 12.4.0 using copies to and from host memory managed by OpenMPI. All processes were assigned to the GPU adjacent to their NUMA region. Each benchmark was run $10$ separate times to demonstrate reproducibility. All plots contain error bars showing the run to run variation of the algorithms.
Figure 9: Multi-lane GPU-aware performance of $2^{27}$ floats on Delta, varying node counts
Figures 8 and 9 compare the performance of a standard MPI_Allreduce operation against the multi-lane optimization. These tests are using a single process per GPU, and using all four GPUs available per node on Delta. While the standard algorithm outperforms multi-lane at the smallest data sizes, the multi-lane algorithm yields significant speedups for larger all-reduces. The speedups for the multi-lane algorithm increase with node count, achieving 2.12.1x speedups at $32$ nodes.
Figure 10: A standard MPI_Allreduce with multiple processes per GPU on Delta
Figure 11: The multi-lane all-reduce with multiple processes per GPU on Delta